OpenStudy (anonymous):

How can I simplify?: ln(1/2)

OpenStudy (anonymous):

\[ \ln(a/b) = \ln(a) -\ln(b) \]

OpenStudy (anonymous):

\[ \ln(1) = 0 \]

OpenStudy (anonymous):

Notice the pattern: \[ \ln(\color{red}a/\color{blue}b) \] \[ \ln(\color{red}1/\color{blue}2) \]

OpenStudy (anonymous):

Oh I remember that okay so I substitute but the calculator gives ln(2) as .6931....

OpenStudy (anonymous):

ln(1) - ln(2) but I don't know what ln(2) is

OpenStudy (anonymous):

You don't need to know, only to simplify. \(\ln(2)\) is the most simple form.

OpenStudy (anonymous):

so the answer would just be ln(1) - ln (2) ?

OpenStudy (anonymous):

Nope, you can simplify \(\ln(1)\)

OpenStudy (anonymous):

See above.

OpenStudy (anonymous):

0 - ln(2) ?

OpenStudy (chrisplusian):

ln(1)=0 so ln(1)-ln(2) is just 0-ln(2) so what @wio is saying is the simplest form is just -ln(2)

OpenStudy (chrisplusian):

yes

OpenStudy (anonymous):

ah thank you :-) @chrisplusian

OpenStudy (chrisplusian):

;) your welcome